# Vistoria: A Multimodal System to Support Fictional Story Writing through Instrumental Text–Image Co-Editing

### v2

@@ 2.3. Instrumental Interaction and Structural Mapping Theory
Instrumental Interaction is central to understanding how users control and refine digital systems. Beaudouin-Lafon (Beaudouin-Lafon, 2000) proposed it as a shift from designing static interface elements to designing instruments that mediate between users and domain objects. A key principle is reification, which transforms abstract commands into persistent, manipulable objects (Kent, 2000). In computing, this elevates implicit system descriptions into explicit first-class entities. In LLM-assisted workflows, this principle appears in modular prompt blocks for structured edits (Yen et al., 2023) and in Textoshop’s reification of abstract text editing commands (e.g., tone adjustment, boolean operations, layers) into direct manipulation tools for text (Masson et al., 2024a). A second principle is polymorphism, where the same instrument applies across contexts (Masson, 2023). This reduces cognitive load by enabling predictable, transferable patterns, e.g., copy–paste works consistently across text, images, and files (Angelini et al., 2015), and scrollbars operate similarly across documents, spreadsheets, and browsers (Masson, 2023). Finally, reuse allows users to replay or adapt prior operations, from macros to redo commands (Riche et al., 2025). Systems like Spacetime exemplify this by objectifying space, time, and actions into persistent containers, enabling edits to be carried forward as manipulable entities (Xia et al., 2018). Together, these principles reduce cognitive burden by externalizing interaction histories, making them manipulable, transferable, and extensible.
In our system, we extend this perspective to AI support for fictional story writing, where imagination must be continuously externalized and revised. We design interaction instruments that reify narrative elements into manipulable objects and, through polymorphism, let the same tools operate across LLM-generated text and images. This integration of reification and polymorphism enables writers to fluidly adjust multimodal outputs as persistent, tangible entities.
Structural Mapping Theory (SMT) is central to understanding how users interpret and compare complex representations. Gentner (Gentner, 1983) proposed SMT to describe how people form abstractions through analogical alignment of relational structures. Rather than relying on surface similarities, SMT highlights the role of alignable differences, which reveal systematic variations and help learners construct new knowledge. Building on this foundation, SMT-informed design emphasizes preserving object structure rather than abstracting it away. At the collection level, showing complete objects allows users to perceive relationships across the dataset (Yan et al., 2022; Wu et al., 2020). At the part level, systems can highlight structural correspondences by surfacing identity matches (literal overlaps) (Glassman et al., 2015), property matches (shared attributes) (Jasim et al., 2022), and conceptual matches (semantic commonalities) (Glassman et al., 2018). These correspondences support the iterative process of aligning, re-aligning, and refining mental models. Prior visualization and learning systems demonstrate the benefits of making such correspondences explicit, between code solutions (Glassman et al., 2015), textual variants (Glassman et al., 2018), or large language model outputs (Arawjo et al., 2024; Gero et al., 2024), helping users notice critical differences, generalize across examples, and avoid disengagement when facing “walls of text.” SMT thus serves as both a cognitive theory and a design constant for interactive systems that deepen users’ reasoning about collections of related objects.
In our system, SMT informs the design of synchronized text–image co-editing: textual fragments are explicitly mapped to visual elements based on their shared properties, matched concepts, or references to the same object, so edits in one modality lead to structurally aligned updates in the other, allowing users to compare, reorganize, and refine story elements across modalities.
@@ 3. Formative Study
Previous research shows that multimodal tools enhance fictional story writing by making abstract concepts tangible, reducing cognitive load, and improving creativity and coherence (Chung et al., 2022; Chung and Kreminski, 2024; Dang et al., 2023; Zhao et al., 2025). However, current tools treat images and texts as supplementary rather than integral to the creative process, leaving unclear how writers actually integrate multiple content types.
To address this gap, we conducted a Wizard-of-Oz co-design study (Dahlbäck et al., 1993) examining how creators use multimodal content (images, text, sketches) when planning and drafting fictional stories (Zhao et al., 2025; Gero et al., 2022; Flower and Hayes, 1981). Our investigation focused on three questions: (1) Multimodal information use: what types of multimodal content users employ and how they leverage these materials for idea generation; (2) Iteration and integration: how creators refine and combine multimodal artifacts in world-building and narrative development; and (3) Organization of inspirations: how creators organize, connect, and refine dispersed inspirations through multimodal manipulation. The WoZ setup simulated AI-assisted visual and textual support while sustaining the impression of an intelligent, interactive system.
@@ 3.1. Process
We designed a Wizard-of-Oz (WoZ) co-design study, positioning participants as active co-designers and treating text, sketches, and images as shared design materials (Dahlbäck et al., 1993; Sanders and Stappers, 2008; Vaajakallio and Mattelmäki, 2014).
@@ 3.1.1. Participants
We recruited 10 participants, each with an two or more years of creative writing experience. The group included three fictional story writers, three animation scriptwriters, two visual film creators, one new media creator, and one online fiction writer. Eight participants held a master’s degree or higher, and two held a bachelor’s degree.
Experimental Setup Three days prior to the session, participants were instructed to prepare a brief fictional story outline consisting of several sentences that followed one of the narrative structures from The Seven Basic Plots (Booker, 2004), which served as the foundation for subsequent ideation and content development. The 90-minute main session took place in Figma(Figma, 2025), which provided a collaborative canvas for free sketching, arranging text–image materials, and treating multimodal inputs as manipulable artifacts.
Session Process During the session, participants engaged in fictional story co-design through activities including generative prompts, collage, and storyboard-like arrangement while interacting with a simulated AI assistant operated by the researcher (wizard). Rather than working toward a fixed output, participants iteratively developed stories of about 300 words while envisioning how the tool itself should behave. They freely requested content through voice, text, or hand-drawn sketches, and called upon the assistant for image and text editing as new needs emerged. The wizard responded in situ to simulate intelligent system behaviors, encouraging naturalistic exploration. To support this process, the experimenter used Claude (Anthropic, 2025) for text generation (drawing on outlines and randomness to stimulate ideation) and Midjourney (Midjourney, 2025) and GPT-4o (OpenAI, 2024) for image generation from prompts, sketches, and editing instructions. Each session concluded with a 30-minute semi-structured interview probing how multimodal materials mediated co-creation, and what interaction patterns and workflows participants desired.
@@ 3.2.1. Using multimodal input to reify vague ideas
As shown in the Table LABEL:tab:formative_grouped, participants utilized multimodal expressions, including text, sketches, and images, as co-design materials to articulate and negotiate intentions with the wizarded system.
@@
Sketches externalized vague intentions and spatial imagination. For example, P4 envisioned a scene where clown JoJo appeared on stage and created a sketch showing “JoJo emerging from a giant spinning wheel” with textual annotations describing the intended atmosphere, hoping AI could elaborate narrative details. Text functioned as the primary medium for conveying intent, allowing participants to express connections between desired content and existing stories while clearly communicating their intentions. Participants frequently used textual annotations to specify story parts or image types for AI generation and guide generation direction. Images communicated style and mood expectations. P3 altered an AI-generated picture’s style by supplying a reference image, while P1 noted, “if possible, I want to use a ’supporting image’–a vague reference picture—as a basis, expecting the AI to generate more detailed images derived from it.”
Table 1. Results of user strategies for manipulating multimodal elements in the WOz co-design study.
Stage Observed behavior User need / Insight N Interaction Example
Multi-modal input Text Generated prose should match the established world while allowing the injection of new elements 83 [Uncaptioned image] [Uncaptioned image]
Multi-modal input Sketches Make spatial relations/composition concrete 8 [Uncaptioned image] [Uncaptioned image]
Planning LLM-Generated images Precisely locate inspiration from images 90 [Uncaptioned image] [Uncaptioned image]
Planning LLM-Generated text Filter usable bits from many generations and reuse them 90 [Uncaptioned image] [Uncaptioned image]
Planning/Translating Text notes / annotations Externalize temporary ideas for re-generation or final writing 30 [Uncaptioned image] [Uncaptioned image]
Planning/Translating Collage elements Recompose fragments from cross-image to form new scenes 12 [Uncaptioned image] [Uncaptioned image]
Planning/Translating Link elements Structure relationships between character / scene / object to connect the plot 40 [Uncaptioned image] [Uncaptioned image]
Translating Reconfigure elements in canvas Reorder scattered ideas by role / time / space into place for global understanding 4 [Uncaptioned image] [Uncaptioned image]
Translating Text (integration) Consolidate AI-generated text and image-inspired content into the draft 10 [Uncaptioned image] [Uncaptioned image]
3.2.2. Text–Image Interplay as Complementary Design Moves
Participants most often sought AI elaboration on characters, objects, and scenes, expressing the need for assistance with character design refinement, setting depictions, or object visualization. Combining image outputs with textual descriptions helped participants enrich their limited knowledge. For example, P7 requested AI-generated designs of an ancient Chinese poison bottle as a narrative element, noting her vague understanding of the concept.
@@ 3.2.2. Text–Image Interplay as Complementary Design Moves
These findings highlight the value of systems that accept heterogeneous multimodal input and help co-designers transform nascent ideas into concrete narrative materials.
@@
Participants perceived text and images as distinct yet complementary in their creative writing. We observed an oscillation between abstraction (text) and concreteness (images) as a recurring co-design pattern, which reveals design opportunities for multimodal authoring tools.
@@
Text as open imagination. Participants described text as a “blank canvas” for associative and structural thinking. P4 noted, “text allows me to imagine many things in my mind,” and P1 emphasized its “infinite imagination on a blank page.” P8 highlighted that text helped set up narrative structure before layering in visuals. This suggests systems should treat text as a flexible space for ideation and intent communication, where ambiguity can be preserved rather than prematurely resolved.
@@
Images as concreteness, inspiration, and feedback. Images grounded abstract ideas, while their randomness often sparked unexpected inspiration. P1 explained: “The randomness in AI-generated images goes beyond what I want or can express; it helps me imagine the next step of the story.” Similarly, P4 refined Lily’s behavior based on an unexpected visual detail, and P3 used images as feedback for progressive refinement. P10 likened it to a “look-and-write exercise” that scaffolds scene construction. These accounts highlight the value of image outputs not just as illustrations but as provocations. Designers can leverage it through image-based iteration, selection, and reinterpretation.
@@
Complementary interplay. Participants emphasized that neither modality sufficed alone: images “set the vibe,” while text reframed meaning. P8 noted, “Images can serve as references for appearance when I don’t have many ideas, while text quickly triggers associations.” We observed participants iteratively moving between text for open-ended imagination and images for concrete grounding, forming a cycle of divergence and convergence. This interplay suggests design opportunities for systems that orchestrate smooth transitions between modalities, enabling users to fluidly oscillate between abstract exploration and concrete elaboration.
@@ 3.2.3. Direct Manipulation of Multimodal Artifacts
Participants expressed a strong interest in treating text and images as manipulable, recombinable design materials. Two recurring practices pointed to design needs for more fluid multimodal manipulation.
Collaging and Recombination. Participants frequently merged elements across outputs to spark new ideas. P9 envisioned combining “the house from the first AI-generated images with the street from the second picture” to construct scenes, while P7 highlighted that “randomly combining characters and scenes” could inspire unexpected connections when accompanied by textual descriptions. As she explained, “if I can see an AI-generated image of my protagonist in one of the scenes, it helps me better imagine potential connections between elements that might otherwise seem unrelated.” Such practices illustrate the potential of collage and recombination as creative strategies.
Granular Editing and Annotation. Beyond recombination, participants desired fine-grained control over outputs. Sticky notes captured details for iteration and served as prompts for later development. Participants also wanted more localized operations, such as regenerating specific regions (P1, P3), extracting and reusing circled image elements (P5), or annotating character personas for refinement (P2). They left narrative prompts for later translation, e.g., P2’s note “Ending could be related to why this postman job even exists.” These behaviors emphasize annotation as both a vehicle for iteration and a bridge to subsequent writing.
Together, these findings suggest systems should enable flexible recombination, localized editing, and traceable annotations to help creators turn multimodal inspiration into actionable writing material.
@@ 3.2.4. From Fragmented Inspirations to Coherent Storylines
While participants often highlighted text or circled inspiring image details, organizing these dispersed fragments into coherent narratives was a persistent challenge. As P2 noted, “everything quickly became too messy on the canvas”, and P4 likened fragments on the canvas to “many cards that required connections,” where narrative coherence depends on linking passages, characters, and settings from scattered parts. Furthermore, P6 wished for mind map–like tools to scaffold this process. Participants also requested clustering notes, surfacing latent relations (e.g., by character/object/setting), and consolidating materials into reusable “setting cards” for ensuring cross-chapter consistency and avoiding logic conflicts (P4, P7).
These challenges point to opportunities for systems that transform fragmented inspirations into structured storylines by supporting clustering, relation mapping, and the creation of reusable narrative units that preserve coherence across iterations.
@@ 3.3. Design Goals
  • DG1: Reifying Multimodal Intentions into Manipulable Artifacts. Informed by findings in Section 3.2.1, our system should reify multimodal inputs (text, sketches, images) into first-class artifacts that can be annotated, rearranged, and recombined. Drawing on Instrumental Interaction, such reification should preserve fleeting intentions as reusable materials for later iteration; following SMT, these artifacts should function as alignable units that support comparison, linkage, and cross-modal reorganization.
  • DG2: Aligning Text and Images for Iterative Creative Exploration. Informed by findings in Section 3.2.2, our system should enable fluid cross-modal iteration: textual edits can be re-visualized, image refinements can inform descriptions, and alternative versions should make differences explicit. Grounded in Dual Coding Theory and SMT, this alignment should support cycles of divergence and convergence by allowing users to oscillate between abstraction (text) and concreteness (images).
  • DG3: Enabling Polymorphic Cross-Modal Manipulation. Informed by findings in Section 3.2.3, our system should support direct manipulation interactions that apply uniformly across modalities, guided by Instrumental Interaction’s principle of polymorphism. This uniformity should reduce switching costs and empower writers to fluidly manipulate and edit textual and visual fragments while maintaining narrative coherence.
  • DG4: Organizing and Reusing Fragments into Coherent Narratives. Informed by findings in Section 3.2.4, our system should cluster and organize fragments and fleeting ideations, surface hidden connections, and consolidate dispersed inspirations into coherent, evolving narratives. It should further support the reuse of narrative units (e.g., characters, settings, objects) to ensure consistency and foster pathways for story progression.

### v3

@@ 2.3. Instrumental Interaction
Instrumental Interaction is central to understanding how users control and refine digital systems. Beaudouin-Lafon (Beaudouin-Lafon, 2000) proposed it as a shift from designing static interface elements to designing instruments that mediate between users and domain objects. A key principle is reification, which transforms abstract commands into persistent, manipulable objects (Kent, 2000). In computing, this elevates implicit system descriptions into explicit first-class entities. In LLM-assisted workflows, this principle appears in modular prompt blocks for structured edits (Yen et al., 2024) and in Textoshop’s reification of abstract image editing commands (e.g., tone adjustment, boolean operations, layers) into direct manipulation tools for text (Masson et al., 2025a). A second principle is polymorphism, where the same instrument applies across contexts (Masson, 2023). This reduces cognitive load by enabling predictable, transferable patterns, e.g., copy–paste works consistently across text, images, and files (Angelini et al., 2015), and scrollbars operate similarly across documents, spreadsheets, and browsers (Masson, 2023). Finally, reuse allows users to replay or adapt prior operations, from macros to redo commands (Riche et al., 2025). Systems like Spacetime exemplify this by objectifying space, time, and actions into persistent containers, enabling edits to be carried forward as manipulable entities (Xia et al., 2018). Together, these principles reduce cognitive burden by externalizing interaction histories, making them manipulable, transferable, and extensible.
In our system, we extend this perspective to the design of multimodal tools for fictional story writing. We reify narrative development as a set of instrumental operations spanning text and imagery. Through polymorphism, the same operation can be applied to both LLM-generated text and images. This combination of reification and polymorphism enables writers to shape multimodal outputs fluidly, aligning with both verbal and nonverbal perceptions.
@@ 3. Formative Study
Previous research shows that multimodal tools enhance fictional story writing by making abstract concepts tangible, reducing cognitive load, and improving creativity and coherence (Chung et al., 2022; Chung and Kreminski, 2024; Dang et al., 2023; Zhao et al., 2025). However, current tools treat images as supplementary rather than integral to the creative process, leaving unclear how writers actually integrate multiple content types into a cohesive workflow.
To address this gap, we conducted a Wizard-of-Oz (WoZ) co-design study (Dahlbäck et al., 1993) examining how creators use multimodal content (images, text, sketches) when planning and drafting fictional stories (Zhao et al., 2025; Gero et al., 2022; Flower and Hayes, 1981). Our investigation focused on three questions: (1) Multimodal information use: what types of multimodal content users employ and how they leverage these materials for idea generation; (2) Iteration and integration: how creators refine and combine multimodal artifacts in world-building and narrative development; and (3) Organization of inspirations: how creators organize, connect, and refine dispersed inspirations through multimodal manipulation. The WoZ setup simulated AI-assisted visual and textual support while sustaining the impression of an intelligent, interactive system.
Our system is designed for writers with intermediate to expert writing expertise, rather than novices who are still learning basic narrative composition. This target group typically possesses established writing habits and a solid understanding of narrative structure. As contemporary writers increasingly incorporate LLMs into their creative workflows (idea generation, style adjustment, etc.) (Chakrabarty et al., 2024b; Lee et al., 2024), we target writers who have hands-on experience using LLMs to assist their writing, even though they may not be experts in multimodal interaction or prompting.
@@ 3.1. Process
We designed a WoZ co-design study, positioning participants as active co-designers and treating text, sketches, and images as shared design materials (Dahlbäck et al., 1993; Sanders and Stappers, 2008; Vaajakallio and Mattelmäki, 2014).
@@ 3.1.1. Participants
For the formative study, we recruited 10 participants through student organizations by sharing our study announcement in group chats, each with at least two years of experience in creative writing. The group included three fictional story writers, three animation scriptwriters, two visual film creators, one new media creator, and one online fiction writer. Eight participants held a master’s degree or higher, and two held a bachelor’s degree. All participants had experience using LLMs to assist in their writing, AI familiarity ranging from casual use (fewer than two days per week, n=5) to daily workflow integration (five or more days per week, n=5).
Experimental Setup. Three days before the session, participants were instructed to prepare a brief fictional story outline consisting of several sentences that followed one of the narrative structures from The Seven Basic Plots (Booker, 2004), which served as the foundation for subsequent ideation and content development. Each session concluded with a 30-minute semi-structured interview probing how multimodal materials mediated co-creation, and what interaction patterns and workflows participants desired.
3.1.2. Wizard-of-Oz System and Session Process
For the WoZ interface, we utilized the canvas in either Figma (31) or Miro (Miro, 2025) as a collaborative space, which was divided into (1) a user-facing “Text Editor” where participants can put in the outline and edit the story, (2) the “Canvas” where generated images and text, and participants’ notes were, and (3) a hidden “Wizard Control Center” (as shown in Figure 2). Participants communicated via voice, text (sticky notes), or hand-drawn sketches while two researchers acted as the “Wizards (system backend)” in real-time to generate outputs, ensuring responsiveness, copying user inputs into separate windows. Researchers ran Claude for text generation, ChatGPT (GPT-4o) and Midjourney [^1] All models were accessed via their commercial web interfaces: https://claude.ai (Anthropic, 2024), https://chatgpt.com (38), and https://www.midjourney.com/ (61) respectively, in June 2025., for simulating the visual engine, then pasting the results back onto the “Canvas.” To ensure consistency, the Wizards followed: (1) input the user’s sketch/text as a literal prompt. Use the initial outline as contextual information for prompts; (2) do not offer creative suggestions unless explicitly asked; (3) to ensure diversity of output style, one researcher generated both texts and images via ChatGPT, and the other researcher generated images via Midjourney and texts via Claude. The two wizards ensured that participants were provided with results that were both timely and diverse.
Figure 2. WoZ System and Study Process (the example of P4) (a) The user edits text in the Text editor; (b) The user writes, sketches, and speaks out their intents; (c)-(d) both wizards paste in the user’s inputs into AI software windows; (e) wizards pasted the generated results back onto the canvas.
The figure depicts the Wizard-of-Oz study setup and interaction process. It shows a text editor where the participant edits story text, a canvas where the participant writes notes, sketches, and expresses intentions, and a hidden wizard control area where researchers input user content into AI tools. The generated images and text are then returned to the canvas, illustrating the end-to-end co-design workflow.
Session Process. During the session, participants engaged in fictional story co-design through activities including generative prompts, collage, and storyboard-like arrangement while interacting with the wizards. Rather than working toward a fixed output, participants iteratively developed stories of about 300 words while envisioning how the tool itself should behave.
@@ 3.2.1. Using Multimodal Input to Reify Vague Ideas
As shown in the Appendix A.1, participants utilized multimodal expressions, including text, sketches, and images, as co-design materials to articulate and negotiate intentions with the wizarded system.
@@
Sketches externalized vague intentions and spatial imagination. For example, P4 envisioned a scene where clown Joko appeared on stage and created a sketch with textual annotations describing the intended atmosphere, hoping AI could elaborate narrative details. Sketches were also frequently used to express ideas that participants found difficult to convey through prompts alone (P6, P7, P8), as they were seen as carrying richer layout and spatial information. Text functioned as the primary medium for conveying intent, allowing participants to express connections between desired content and existing stories. Participants also frequently used textual annotations to specify story parts or image types for AI-generated content and to guide the direction of content generation. They also used textual annotations to record the inspirations they received and how those ideas might be used in writing. Images communicated style and mood expectations. P3 altered an AI-generated picture’s style by supplying a reference image, while P1 noted, “If possible, I want to use a ‘supporting image’—a vague reference picture—as a basis, expecting the AI to generate more detailed images derived from it.” In addition, sketches were often paired with images or text to further articulate the intentions participants held in mind (P4, P7).
@@
These findings highlighted the value of systems that accept different modality input and help co-designers transform nascent ideas into concrete narrative materials.
3.2.2. Image-Text Interplay as Complementary Design Moves
@@
Participants perceived text and images as distinct yet complementary, and we observed them often switching between abstraction (text) and concreteness (images) as a recurring co-design pattern.
@@
Text as open imagination. Participants described text as a “blank canvas” for boundless imagination. P4 noted, “Text allows me to imagine many things in my mind,” and P1 emphasized text as “infinite imagination on a blank page.” P8 highlighted that text helped set up the narrative structure before layering in visuals. This suggests systems should treat text as a flexible space for ideation and intent communication, where ambiguity can be preserved rather than early resolved.
@@
Images as concreteness, inspiration, and feedback. Images grounded abstract ideas, while their randomness often sparked unexpected inspiration. P1 explained: “The randomness in AI-generated images goes beyond what I want or can express; it helps me imagine the next step of the story.” Similarly, P10 regarded referring to images as a “look-and-write exercise” that scaffolds scene construction. These accounts highlight the value of image outputs not just as illustrations but as provocations.
@@
Complementary interplay. Participants emphasized that neither modality sufficed alone: images “set the vibe,” while text reframed meaning. P8 noted, “Images can serve as references for appearance when I don’t have many ideas, while text quickly triggers associations.” We observed participants iteratively moving between text for open-ended imagination and images for concrete grounding, forming a cycle of divergence and convergence (P4, P5, P8, P9, P10). Four participants (P1, P2, P5, P7) also expressed a desire for text and image changes to be synchronized, so they would not need to constantly cross-check and compare updates across modalities. In addition, the concurrent presentation of text and images further facilitates narrative expression. As P7 explained, “ When text and images appear together, I can describe the scene while referencing the image and directly adopt AI-generated text when it proves useful. This also allows inspiration to emerge.”
This interplay suggested design opportunities for systems that incorporated smooth transitions between text and images while aligning these two modalities, enabling participants to fluidly move between abstract exploration and concrete elaboration
@@ 3.2.3. Direct Manipulation of Multimodal Artifacts
Participants expressed a strong interest in treating text and images as manipulable, recombinable design materials. Two recurring practices pointed to design needs for more direct multimodal manipulation.
Collaging and Recombination. Participants frequently merged elements across outputs to spark new ideas. P9 envisioned combining “the house from the first AI-generated images with the street from the second picture” to construct scenes, while P7 highlighted that “randomly combining characters and scenes” could inspire unexpected connections when accompanied by textual descriptions. She also expressed that visualizing hard-to-imagine scenes makes it easier to write richer narrative descriptions. Such practices illustrated the potential of collage and recombination as creative strategies.
Granular Editing and Annotation. Beyond recombination, participants desired fine-grained control over outputs. They use sticky notes to capture details for iteration and to serve as prompts for later development. Participants also wanted more localized operations, such as regenerating specific regions (P1, P3), extracting and reusing circled image elements (P5), or annotating character personas for refinement (P2). These behaviors emphasized editing and annotation as both vehicles for iteration and a bridge to subsequent writing.
Together, these findings suggested systems should enable flexible recombination, localized editing, and traceable annotations to help creators iteratively refine narrative materials.
@@ 3.2.4. From Fragmented Inspirations to Coherent Storylines
While participants often highlighted text or circled inspiring image details, organizing these dispersed fragments into coherent narratives was a persistent challenge in the planning phase. As P2 noted, “Everything quickly became too messy on the canvas,” and P4 likened fragments on the canvas to “many cards that required connections,” where narrative coherence depends on linking passages, characters, and settings from scattered parts. Furthermore, P6 wished for mind map–like tools to scaffold this process. Participants also requested clustering notes, surfacing latent relations (e.g., by character/object/setting), and consolidating materials into reusable “setting cards” to ensure cross-chapter consistency and avoid logic conflicts (P4, P7).
These challenges pointed to opportunities for systems that transform fragmented inspirations into structured storylines by supporting clustering, relation mapping, and the creation of reusable narrative units that preserve coherence across iterations.
@@ 3.3. Design Goals
Figure 3. Vistoria’s interface: a left Text Editor, a central Cluster panel (can be collapsed when not used), and a right free-form Canvas.
The figure presents Vistoria’s interface layout. It consists of a left-side Text Editor displaying the current story draft, a central Cluster panel that aggregates highlighted narrative elements, and a right-side free-form Canvas for sketching, placing images, writing text, and performing image–text co-editing operations.
  • DG1: Supporting the Expression of Ideas through Multimodality. Grounded in the findings in Section 3.2.1, our system should provide multimodal mechanisms combining sketches, text, and images to capture early intention, imprecise expressions, and help transform them into concrete narrative materials for further iterative editing or re-organizing.
  • DG2: Aligning Text and Images for Iterative Creative Exploration. Informed by findings in Section 3.2.2, our system should enable fluid cross-modal iteration: textual edits can be re-visualized, and image refinements can inform text descriptions. Grounded in Dual Coding Theory, text and image updates should also be synchronized to more effectively align verbal and nonverbal perception.
  • DG3: Enabling Polymorphic Cross-Modal Manipulation. Informed by findings in Section 3.2.3, our system should support direct manipulation interactions for both text and images. Guided by Instrumental Interaction’s principle of polymorphism, we should design the same instrument for cross-modal editing to reduce switching costs and enable writers to manipulate textual and visual fragments while maintaining narrative coherence.
  • DG4: Organizing and Reusing Fragments into Coherent Narratives. Informed by the findings in Section 3.2.4, our system should support clustering and organizing fragments and fleeting ideations during the exploratory phase, surface latent connections, and consolidate dispersed inspirations into coherent, evolving narrative structures that support translation into final writing.